You can afford a $450 per month car payment. You've found a 4 year loan at 3% interest. How big of a loan can you afford?

Answer by math_tutor2020(3816)


Monthly payment formula
P = (L*i)/(1 - (1+i)^(-n))
where,
P = monthly payment
L = loan amount
i = interest rate per month
n = number of months

When solving for variable L we get
L = P*(1 - (1+i)^(-n))/i

In this case we have these inputs
P = 450
i = 0.03/12 = 0.0025
n = 12*4 = 48

So,
L = P*(1 - (1+i)^(-n))/i
L = 450*(1 - (1+0.0025)^(-48))/0.0025
L = 20330.41258572
L = 20330.41 is the most you can borrow.